The lateral edge of a pyramidal church spire is 61 ft. Each side of the

octagonal base is 22 ft. What will be the cost of painting the spire at

$0.025 a square foot?



1
Expert's answer
2021-11-30T02:40:13-0500

Lateral area "=\dfrac{1}{2}" Perimeter of the base "\times" Slant height

Lateral area "=\dfrac{1}{2}(8\cdot22)\times61=5368 ({ft}^2)"


"Cost=\$0.025/{ft}^2\cdot5368{ft}^2=\$134.20"

The cost of painting the spire is $134.20.